# Graduate Student Member, IEEE

### v1

@@ 3.1 Acquisition Results from IQ Snapshots Collected in the Lunar Domain
Fig. 6(a) and Fig. 6(b) present the distribution of $C/N_{0}$ values for each signal on the L1 and L5 bands, together with the number of successful acquisitions per signal. The analysis indicates that expanding the set of processed signals to include those of additional GNSS, RNSS, and SBAS constellations substantially increases the total number of successful acquisitions. On the L1 band, the largest contributions arise from SBAS and BeiDou, whereas on the L5 band, BeiDou and IRNSS are the most significant. Notably, the NavIC L5 signal characterized by a chip rate of 1.023 Mchip/s, which is 10 times lower with respect to chip rate of all the other L5 signals, so it is uniquely oversampled. It is worth recalling that all L5 SC operations were instead performed at a nominal sampling rate of 24 MHz.
@@
Figure 6: Distribution of the carrier-to-noise density ratio ($C/N_{0}$) for different GNSS signals in the L1 band. For each signal, the central line of the box represents the median $C/N_{0}$, the box spans the interquartile range (25th–75th percentiles), and the whiskers extend to the most extreme values excluding outliers. Outliers, when present, are indicated by individual markers. The lower panel reports the number of occurrences for each signal.
@@ 3.2 Simulated Signal Availability
For all subsequent analyses, a threshold of 26 dB-Hz is used. This value provides a conservative margin, since the LuGRE receiver has successfully tracked signals with an estimated $C/N_{0}$ of about 24 dB-Hz [JournalDiMissione]. The selected threshold also improves robustness against possible modeling inaccuracies and other sources of signal degradation. It is important to note that the results shown here, besides the selected sensitivity threshold, are receiver-agnostic; different receivers may adopt different acquisition and tracking strategies, and their effective thresholds can vary depending on both environmental conditions and internal receiver state.
Fig. 7(a) and Fig. 7(b) illustrate the percentage of the nominal RTP operation duration during which $n$ signals from $n$ distinct satellites exceed the selected $C/N_{0}$ threshold. Specifically, Fig. 7(a) presents the multi-frequency, multi-constellation scenario including only GPS and Galileo, while Fig. 7(b) shows the case where an additional GNSS, RNSS, and SBAS constellations are also considered. The figures show an increase in the availability introduced by the additional constellations; anyway, some operations are still characterized by a reduced number of satellites in radiometric visibility, such as the case of OP74_0 and OP27_0. Including BeiDou, as well as the RNSS and SBAS constellations, increases the percentage of epochs with at least four satellites in radiometric visibility from 11.7% to 46.1%. The adopted threshold, i.e., $n\geq 4$, was considered as it represents the minimum requirement for autonomous receiver state estimation, assuming the use of limited a priori information to guarantee solution convergence (e.g. through a Least-Squares estimation). This condition allows for the reliable initialization of more advanced orbit determination filters such as Bayesian estimation filters devoted to Precise orbit determination (POD).
@@ 3.3 Dilution of precision
A partial reduction in Geometric Dilution Of Precision (GDOP) resulting from the inclusion of additional constellations can be observed in Fig. 8, which shows the GDOP values for all the operations in which the evaluation was possible. As it can be noticed, a wide range of GDOP values is observed, reaching a maximum at almost 10,000 and a minimum close to 300. During some operations, including SBAS and RNSS satellites in the set of satellites used for GDOP computation leads to an improvement in GDOP. This effect is particularly evident in operations OP17_0, OP23_0, OP37_0, OP38_0 and OP76_0. When GDOP cannot be evaluated without SBAS or RNSS satellites, the contribution to the availability increase becomes evident. Fig. 9 illustrates a distinctive operational condition, showing the $C/N_{0}$ values generated by the extended SSV simulator. During OP38_0, the Galileo E1B signal exhibits a consistent pattern, as reported in [JournalDiMissione]; signals from SVs on orbital plane B progressively fade, while those on plane A remain visible with higher $C/N_{0}$ despite increasingly prolonged Earth-shadowing periods. The simulations also show that the Medium Earth Orbit (MEO) SVs of the BeiDou constellation follow a similar fading pattern, whereas the BeiDou IGSO and GEO satellites do not. This naturally mitigate the unfortunate phasing phenomena at lunar surface. However, it is important to remark that the use of SBAS signals for ranging purposes may require more accurate ephemerides and additional corrections with respect to those broadcast by the SVs themselves [kahr2016analysis].
@@
Figure 7: Multi-constellation, multi-frequency availability as the percentage of nominal RTP operation with $n$ signals from $n$ distinct GNSS satellites above the $C/N_{0}$ threshold, shown for $n\geq 4$.
@@ 4 Conclusions
The experimental results are used to calibrate the extended SSV simulator. The simulation of mission operations shows that extending the analysis beyond GPS and Galileo to include BeiDou, regional navigation satellite systems such as IRNSS (NavIC) and QZSS, as well as SBAS constellations, leads to a significant improvement in signal availability. In particular, in a dual-frequency multi-constellation configuration, the percentage of total epochs in which at least four satellites are in radiometric visibility increases from 11% to 46% with respect to the GPS- and Galileo-only case. Nevertheless, despite this improvement, there remain mission phases in which fewer than four satellites are simultaneously in radiometric visibility. In addition to improving availability, processing multiple constellations increases the probability of meeting the minimum visibility required for a priori-free convergent navigation solutions, allowing on-demand filter initialization without long delays, although high-sensitivity acquisition may still be required. The inclusion of RNSS and SBAS GEO constellations further enhances orbital geometry diversity, mitigating geometry-dependent effects such as orbital-plane–related signal fading observed during specific operational periods (e.g., OP38_0), and thereby improving availability, at the cost of requiring additional corrections for the SBAS SVs [kahr2016analysis].
@@ Acknowledgment
{IEEEbiography}
[[Uncaptioned image]]Lorenzo Sciacca (Graduate Student Member, IEEE) received his M.Sc. in Communications and Computer Network Engineering from Politecnico di Torino, Italy, in 2024. He is currently pursuing a Ph.D. at the Department of Electronics and Telecommunications, Politecnico di Torino, as part of the Navigation Signal Analysis and Simulation (NavSAS) group. His research focuses on developing high-sensitivity, environment-aware GNSS receivers to improve navigation in challenging conditions, including lunar and cis-lunar environments. He has been involved in the Lunar GNSS Receiver Experiment (LuGRE) since 2025.
{IEEEbiography}
[[Uncaptioned image]]Alex Minetto (Member, IEEE) received the B.Sc. (2013), M.Sc. (2015), and Ph.D. with honors (2020) in telecommunications and electrical engineering from Politecnico di Torino, where he is now Assistant Professor in the Department of Electronics and Telecommunications and a member of the NavSAS research group. His work focuses on advanced GNSS signal analysis, receiver design, and Bayesian state estimation. He has contributed to ESA and EUSPA projects and since 2021 has been involved in the Lunar GNSS Receiver Experiment (LuGRE) as a mission’s Science Team member.).
{IEEEbiography}
[[Uncaptioned image]]Andrea Nardin (Member, IEEE) received the M.Sc. (2018) and Ph.D. with honors (2023) in telecommunications and electrical engineering from Politecnico di Torino, where he is now Assistant Professor in the Department of Electronics and Telecommunications and a member of the NavSAS research group. In 2021, he was a Visiting Doctoral Researcher at Northeastern University, Boston. His research interests include signal processing architectures and signal design for GNSS and next-generation PNT systems.
{IEEEbiography}
[[Uncaptioned image]]Fabio Dovis (Member, IEEE) received the M.Sc. (1996) and Ph.D. (2000) from Politecnico di Torino, where he is now Full Professor in the Department of Electronics and Telecommunications and coordinator of the NavSAS Research Group. His research focuses on GNSS receiver design, advanced signal processing for interference and multipath mitigation, and ionospheric monitoring. He has extensive experience in internation projects, collaborations with industry and research institutions, and serves on the IEEE AESS Navigation Systems Panel.
{IEEEbiography}
[[Uncaptioned image]]Luca Canzian joined Qascom in 2015 and he is currently leading the R&D domain area. He has worked in several projects with ESA, ASI, NASA, the European Commission and Industry. His main expertise involves ground-based and space-based location systems, detection and location of interference signals, navigation using signals of opportunity, lunar navigation, inertial navigation systems, orbit determination techniques, GNSS authentication and anti-spoofing techniques. He holds a MSc degree and a PhD in Electrical Engineering from University of Padova (Italy).
{IEEEbiography}
[[Uncaptioned image]]Mario Musmeci received a degree in physics from the University of Rome “La Sapienza.” He has worked with major Italian space companies, ESA, and the European Commission on the Galileo program, and co-founded a start-up exploiting a patent based on Galileo signals for reliable traceability of “Made in Italy” products. Since 2016, he has been with ASI’s Programs Directorate, serving as technical manager for LuGRE on NASA’s Firefly Blue Ghost mission, demonstrating GNSS reception on the lunar surface.
{IEEEbiography}
[[Uncaptioned image]]Claudia Facchinetti is a researcher and technologist at the Italian Space Agency (ASI), specializing in Earth observation and space navigation technologies. She holds a Ph.D. in Space Science and Technology and has contributed to numerous scientific projects and space missions, including studies on SAR radar and hyperspectral sensors. She is a key member of the LuGRE (Lunar GNSS Receiver Experiment) team, a joint NASA-ASI initiative that successfully demonstrated GNSS-based navigation on the Moon during the Firefly Blue Ghost Mission 1 in 2025. Additionally, she is responsible for developing Earth observation missions at ASI.
{IEEEbiography}
[[Uncaptioned image]]Giancarlo Varacalli received the degree in Electronic Engineering and holds a Master’s degree in Space Systems Engineering from Delft University of Technology. He is currently Head of the Telecommunications and Navigation Department at the Italian Space Agency (ASI). He has held several technical and managerial positions within ASI and serves as Italy’s delegate to ESA programme boards in satellite communications and navigation. His professional interests include satellite telecommunications, navigation systems, and space system engineering.

### v2

@@ 4.1 Acquisition Results from IQ Snapshots Collected in the Lunar Domain
Fig. 6(a) and Fig. 6(b) present the distribution of $C/N_{0}$ values for each signal on the L1 and L5 bands, together with the number of successful acquisitions per signal. The analysis indicates that expanding the set of processed signals to include those of additional GNSS, RNSS, and SBAS constellations substantially increases the total number of successful acquisitions. On the L1 band, the largest contributions arise from SBAS and BeiDou, whereas on the L5 band, BeiDou and IRNSS are the most significant. Notably, the NavIC L5 signal is characterized by a chip rate of 1.023 Mchip/s, which is 10 times lower than the chip rate of all other L5 signals, making it uniquely oversampled. It is worth recalling that all L5 SC operations were instead performed at a nominal sampling rate of 24 MHz.
@@
Figure 6: Distribution of the carrier-to-noise density ratio ($C/N_{0}$) for different GNSS signals in the L1 and L5 bands. For each signal, the central line of the box represents the median $C/N_{0}$, the box spans the interquartile range (25th–75th percentiles), and the whiskers extend to the most extreme values excluding outliers. Outliers, when present, are indicated by individual markers. The lower panel reports the number of occurrences for each signal.
@@ 4.2 Simulated Signal Availability
For all subsequent analyses, a threshold of 26 dB-Hz is used. This value provides a conservative margin, since the LuGRE receiver has successfully tracked signals with an estimated $C/N_{0}$ of about 24 dB-Hz [17]. The selected threshold also improves robustness against possible modeling inaccuracies and other sources of signal degradation. It is important to note that the results shown here, besides the selected sensitivity threshold, are receiver-agnostic; different receivers may adopt different acquisition and tracking strategies, and their effective thresholds can vary depending on both environmental conditions and internal receiver state.
Fig. 7(a) and Fig. 7(b) illustrate the percentage of the nominal RTP operation duration during which $n$ signals from $n$ distinct satellites exceed the selected $C/N_{0}$ threshold. Specifically, Fig. 7(a) presents the multi-frequency, multi-constellation scenario including only GPS and Galileo, while Fig. 7(b) shows the case where additional GNSS, RNSS, and SBAS constellations are also considered. The figures show an increase in the availability introduced by the additional constellations; however, some operations are still characterized by a reduced number of satellites in radiometric visibility, such as the case of OP74_0 and OP27_0. Including BeiDou, as well as the RNSS and SBAS constellations, increases the percentage of epochs with at least four satellites in radiometric visibility from 11.7% to 46.1%. The adopted threshold, i.e., $n\geq 4$, was considered as it represents the minimum requirement for autonomous receiver state estimation, assuming the use of limited a priori information to guarantee solution convergence (e.g. through a Least-Squares estimation). This condition allows for the reliable initialization of more advanced orbit determination filters such as Bayesian estimation filters devoted to Precise orbit determination (POD).
@@ 4.3 Dilution of precision
A partial reduction in Geometric Dilution Of Precision (GDOP) resulting from the inclusion of additional constellations can be observed in Fig. 8, which shows the GDOP values for all the operations in which the evaluation was possible. As it can be noticed, a wide range of GDOP values is observed, reaching a maximum at almost 10,000 and a minimum close to 300. During some operations, including SBAS and RNSS satellites in the set of satellites used for GDOP computation leads to an improvement in GDOP. This effect is particularly evident in operations OP17_0, OP23_0, OP37_0, OP38_0 and OP76_0. When GDOP cannot be evaluated without SBAS or RNSS satellites, the contribution to the availability increase becomes evident. Fig. 9 illustrates a distinctive operational condition, showing the $C/N_{0}$ values generated by the extended SSV simulator. During OP38_0, the Galileo E1B signal exhibits a consistent pattern, as reported in [17]; signals from SVs on orbital plane B progressively fade, while those on plane A remain visible with higher $C/N_{0}$ despite increasingly prolonged Earth-shadowing periods. The simulations also show that the Medium Earth Orbit (MEO) SVs of the BeiDou constellation follow a similar fading pattern, whereas the BeiDou IGSO and GEO satellites do not. This naturally mitigates the unfortunate phasing phenomena at lunar surface. However, it is important to remark that the use of SBAS signals for ranging purposes may require more accurate ephemerides and additional corrections with respect to those broadcast by the SVs themselves [8].
@@
Figure 7: Multi-constellation, multi-frequency availability as the percentage of nominal RTP operation with $n$ signals from $n$ distinct GNSS satellites above the $C/N_{0}$ threshold. Percentages are shown for $n\geq 4$.
@@ 5 Conclusions
The experimental results are used to calibrate the extended SSV simulator. The simulation of mission operations shows that extending the analysis beyond GPS and Galileo to include BeiDou, regional navigation satellite systems such as IRNSS (NavIC) and QZSS, as well as SBAS constellations, leads to a significant improvement in signal availability. In particular, in a dual-frequency multi-constellation configuration, the percentage of total epochs in which at least four satellites are in radiometric visibility increases from 11% to 46% with respect to the GPS- and Galileo-only case. Nevertheless, despite this improvement, there remain mission phases in which fewer than four satellites are simultaneously in radiometric visibility. In addition to improving availability, processing multiple constellations increases the probability of meeting the minimum visibility required for a-priori-free convergent navigation solutions, allowing on-demand filter initialization without long delays, although high-sensitivity acquisition may still be required. The inclusion of RNSS and SBAS GEO constellations further enhances orbital geometry diversity, mitigating geometry-dependent effects such as orbital-plane–related signal fading observed during specific operational periods (e.g., OP38_0), and thereby improving availability, at the cost of requiring additional corrections for the SBAS SVs [8].